OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. - Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation.

OpenStudy (anonymous):

f(x) = 15x + 85 The total price is based on the number of charms bought because for every charm bought, the total price raises by $15.00. for instance : if only 1 charm is bought : f(1) = 15(1) + 85 f(1) = 15 + 85 f(1) = 100 so if 1 charm is bought, the total is $100 and if 2 charms are bought.. f(2) = 15(2) + 85 f(2) = 30 + 85 f(2) = 115 so if 2 charms are bought, the total is $115, which is $15 more....so you see, each charm raises the total price by $15.00.